When positioned at 90° to a magnetic field, a wire of length 1 m carrying a current of 4 A experiences a force of 0.2 N. What is the strength of the magnetic field?

1
Expert's answer
2022-05-22T14:39:13-0400
"F=BIl\sin\theta"

"B=\frac{F}{Il\sin\theta}=\frac{0.2}{4*1*\sin90^\circ}=0.05\:\rm T"